OpenStudy (anonymous):

find the derivative of the function: f(x)= x^2+5x+2/e^x

OpenStudy (baldymcgee6):

What rule should you use?

OpenStudy (anonymous):

There is one division sign(/) there.. So, which rule you will follow ??

OpenStudy (baldymcgee6):

\[f(x)= \frac{ x^2 +5x+2 }{ e^x }\]I assume it is this?

OpenStudy (anonymous):

Yes it is... I don't think there is doubt in this...

OpenStudy (anonymous):

Or may be... Ha ha ha ha....

OpenStudy (baldymcgee6):

there is a difference between f(x)= x^2+5x+2/e^x and f(x)= (x^2+5x+2)/e^x

OpenStudy (australopithecus):

splitting it up g(x) = x^2+5x+2 g'(x) = 2x + 5 c(x) = 1/e^x = e^-x c'(x) = -e^-x sub that into product rule f'(x) = g'(x)c(x) + c'(x)g(x) f'(x) = (2x+5)/e^x + x^2+5x+2/(-e^x)

OpenStudy (australopithecus):

any questions?

OpenStudy (anonymous):

You have given the whole answer.. And who is going to ask questions here.. All have got all answers...

OpenStudy (australopithecus):

I was wondering if the person had any questions about how I came to the solution, yes I provided the answer but I showed my method so that the person my learn from it

OpenStudy (australopithecus):

may*

OpenStudy (anonymous):

What if the asker is only interested in whole solution and not in how it came ??

OpenStudy (australopithecus):

well then he is a fool

OpenStudy (australopithecus):

and a jerk for wasting peoples time as he could easily just use wolframalpha to gain the solution

OpenStudy (baldymcgee6):

@Australopithecus wolfram alpha will also show you the steps, as you did.. some people need an explanation.

OpenStudy (anonymous):

And I think you are greatest fool too if you know that he can use Wolfram Alpha and then also you are giving whole solution in one go... Please don't mind because I am mad..

OpenStudy (australopithecus):

My steps are more intuitive though than wolframalpha

OpenStudy (anonymous):

But these steps are useless if the asker is not interested in how these steps came..

OpenStudy (anonymous):

Leave all this here only.. Yes your steps are more intuitive... @Australopithecus

OpenStudy (australopithecus):

Look waterineyes I can see where you are coming from, I mean not giving the solution to a problem will prevent immature people who are just looking for a solution from posting on the site and not wasting everyones time. In the end it doesn't matter though because these people are always going to show up. You can't force someone to learn

OpenStudy (australopithecus):

Everything we do is frivolous in the end we are but dust in the wind. So don't hassle me with what I post, I have good intentions, if you wanted to lead this person to the answer that is fine I wont butt in, but I didn't see you helping him and I'm done explaining myself to you.

OpenStudy (baldymcgee6):

If people don't just give out answers like you, they will be pressed to learn the material; or better yet, you could help them learn. The policy of the site clearly states to LEAD the asker to a solution, as opposed to giving them all the answers so they can copy it down on to their homework.

OpenStudy (baldymcgee6):

Both me and @waterineyes were here helping the asker before you even showed up. Clearly you DID butt in.

OpenStudy (anonymous):

@baldymcgee6 leave it buddy...

OpenStudy (australopithecus):

Well chances are he wont only be doing one derivative in his life time, so if he wants to just copy the answer fine but it is going to be detrimental to him when he goes to write a test or answer other questions

OpenStudy (baldymcgee6):

@waterineyes I agree, maybe he was right that some people can't learn.

OpenStudy (australopithecus):

I clearly asked if he understood the solution

OpenStudy (anonymous):

@baldymcgee6 let us go to other problems and help others...